BRICS Notes Series, 1995

March 22, 2004

This document is also available as PostScript and DVI.

References

[NS-95-6] Abstract.

Aravind Srinivasan. The Role of Randomness in Computation. November 1995. iv+99 pp.

[NS-95-5] Abstract.

Robert Paige. Analysis and Transformation of Set-Theoretic Languages. Mini-Course. August 1995. iv+157 pp.

[NS-95-4] PostScript.

Yuri Gurevich and Egon Brger. Evolving Algebras. Mini-Course. July 1995. iv+222 pp.

[NS-95-3] Abstract, PostScript, DVI.

Andrew D. Gordon. Bisimilarity as a Theory of Functional Programming. Mini-Course. July 1995. iv+59 pp.

[NS-95-2] Abstract, PostScript.

Uffe H. Engberg, Kim G. Larsen, and Arne Skou, editors. Proceedings of the Workshop on Tools and Algorithms for The Construction and Analysis of Systems, TACAS (Aarhus, Denmark, 19–20 May, 1995), May 1995. vi+334 pp. Selected papers appears in Brinksma, Cleaveland, Larsen, Margaria and Steffen, editors, Tools and Algorithms for The Construction and Analysis of Systems: International Workshop, TACAS ’95 Selected Papers, LNCS 1019, 1995.

[NS-95-1] Abstract, PostScript.

Igor Walukiewicz. Notes on the Propositional μ-calculus: Completeness and Related Results. February 1995. 54 pp.